# Multi-qubit Dynamical Decoupling for Enhanced Crosstalk Suppression

### 08 Mar 2024 v1

@@ Appendix A Various Experimental Configurations
There are two variants of CX gates in IBM quantum hardware: Echoed Cross-Resonance CX (ECR), and Direct CX (DCX). We select several qubit pairs from ibmq_cairo including both ECR and DCX types, and list their respective amplitudes, duration, and error rates in Table 2. Note that, CX gate is natively uni-direction in IBM quantum hardware, and the opposite direction is realized by adding extra single qubit gates. Therefore, we show explicitly the control and target qubits for each qubit pair in the table. Generally, DCX gate exhibits a lower error rate and shorter duration compared with ECR gate. However, DCX gate is more susceptible to crosstalk, as illustrated in Fig. 6. Here, the CNOT error rate, derived from SRB, exhibits a notable increase compared to the rate derived from independent RB for the DCX type of CNOT gate, whereas the CNOT error rate for the ECR CNOT only experiences a slight increase.
@@
We conduct both idle-idle and driven-idle experiments on adjacent qubit pairs across three distinct scenarios: (DCX, DCX), (ECR, DCX), and (ECR, ECR). Given the adjacency of the two pairs, one qubit from the first pair is positioned next to a qubit from the second pair, following different configurations: control-control, control-target, or target-target. The results of these experiments are displayed in Fig. 7. In the case of driven-idle experiments, staggered DD consistently demonstrates superior performance over standard DD. However, for the idle-idle experiments, the performance of staggered DD exhibits a lack of consistency.
@@
Figure 7: Free evolution and crosstalk experiment results on pairs $[(q_{1},q_{0})-(q_{4},q_{7})]$
@@
Figure 7: Free evolution and crosstalk experiment results on pairs $[(q_{1},q_{0})-(q_{4},q_{7})]$
@@
The exploration of circuit fidelity for both idle-idle and driven-idle experiments reveal that the insertion of $X2$ sequence leads to worse outcomes than the baseline, as shown in Fig. 8, which is potentially due to the imperfections in the X gate. However, this was not reproducible in simulations with an error model informed by quantum process tomography. Rather, we suspect an observation similar to Ref. [17] performed on IBM quantum hardware, where fidelity decreases due to the asymmetry of the trajectory around the Bloch sphere during qubit rotations. Since an X gate corresponds to a $\pi$ rotation around the $x$-axis in one direction, this asymmetry accumulates after two consecutive X gates, thereby resulting in poorer performance compared to the baseline where no DD sequence is inserted. On the other hand, the Xp and Xm operations, which correspond to a $\pi$ rotation around the $X$ axis in opposite directions, effectively mitigate this asymmetry.
@@ Appendix C Noisy Simulation Results
The procedure begins by inserting $X2pm$ along with its staggered counterpart into the delay of the Ramsey experiment. Within these DD-inserted circuits, we incorporate an additional $ZZ(\theta)$ gate prior to each delay, with $\theta=\omega t$, serving to simulate $ZZ$ crosstalk across time evolution $t$. We evaluate the discrepancies in detuning frequency when inserting either standard or staggered DD sequences into Ramsey experiment circuits. This comparison is performed under two conditions: with and without the introduction of $ZZ$ crosstalk. The resulting detuning frequencies are detailed in Table 3. Staggered DD is shown to be highly effective in suppressing the impacts of $ZZ$ crosstalk, evident in its ability to maintain a consistent detuning frequency in comparison to the non-crosstalk scenario. Conversely, the presence of $ZZ$ crosstalk instigates a frequency shift when inserting standard DD. This noise simulation outcome aligns with the theoretical analysis, validating the capability of staggered DD in canceling out $ZZ$ crosstalk.

### 14 Mar 2024 v2

@@ Appendix A Various Experimental Configurations
There are two variants of CX gates in IBM quantum hardware: Echoed Cross-Resonance CX (ECR), and Direct CX (DCX). We select several qubit pairs from ibmq_cairo including both ECR and DCX types, and list their respective amplitudes, duration, and error rates in table 2. Note that, CX gate is natively uni-direction in IBM quantum hardware, and the opposite direction is realized by adding extra single qubit gates. Therefore, we show explicitly the control and target qubits for each qubit pair in the table. Generally, DCX gate exhibits a lower error rate and shorter duration compared with ECR gate. However, DCX gate is more susceptible to crosstalk, as illustrated in figure 6. Here, the CNOT error rate, derived from SRB, exhibits a notable increase compared to the error derived from independent RB for DCX, whereas the CNOT error rate for the ECR only experiences a slight increase.
@@
We conduct both idle-idle and driven-idle experiments on adjacent qubit pairs across three distinct scenarios: (DCX, DCX), (ECR, DCX), and (ECR, ECR). Given the adjacency of the two pairs, one qubit from the first pair is positioned next to a qubit from the second pair, following different configurations: control-control, control-target, or target-target. The results of these experiments are displayed in figure 7. In the case of driven-idle experiments, staggered DD consistently demonstrates superior performance over standard DD. However, for the idle-idle experiments, the performance of staggered DD exhibits a lack of consistency.
@@
Figure 7: Idle-idle and driven-idle experiment results on pairs $[(q_{1},q_{0})-(q_{4},q_{7})]$
@@
Figure 7: Idle-idle and driven-idle experiment results on pairs $[(q_{1},q_{0})-(q_{4},q_{7})]$
@@
The exploration of circuit fidelity for both idle-idle and driven-idle experiments reveal that the insertion of $X2$ sequence leads to worse outcomes than the baseline, as shown in figure 8, which is potentially due to the imperfections in the X gate. However, this was not reproducible in simulations with an error model informed by quantum process tomography. Rather, we suspect an observation similar to [17] performed on IBM quantum hardware, where fidelity decreases due to the asymmetry of the trajectory around the Bloch sphere during qubit rotations. Since an X gate corresponds to a $\pi$ rotation around the $x$-axis in one direction, this asymmetry accumulates after two consecutive X gates, thereby resulting in poorer performance compared to the baseline where no DD sequence is inserted. On the other hand, the Xp and Xm operations, which correspond to a $\pi$ rotation around the $X$ axis in opposite directions, effectively mitigate this asymmetry.
@@ Appendix C Noisy Simulation Results
The procedure begins by inserting X2pm along with its staggered counterpart into the delay of the Ramsey experiment. Within these DD-inserted circuits, we incorporate an additional $ZZ(\theta)$ gate prior to each delay, with $\theta=\omega t$, serving to simulate $ZZ$ crosstalk across time evolution $t$. We evaluate the discrepancies in detuning frequency when inserting either standard or staggered DD sequences into Ramsey experiment circuits. This comparison is performed under two conditions: with and without the introduction of $ZZ$ crosstalk. The resulting detuning frequencies are detailed in table 3. Staggered DD is shown to be highly effective in suppressing the impact of $ZZ$ crosstalk, evident in its ability to maintain a consistent detuning frequency in comparison to the non-crosstalk scenario. Conversely, the presence of $ZZ$ crosstalk instigates a frequency shift when inserting standard DD. This noise simulation outcome aligns with the theoretical analysis, validating the capability of staggered DD in canceling out $ZZ$ crosstalk.